A 2500 kg car is moving down the highway at a velocity of 30.5 m/s when it bumps into the car ahead of it at which has a mass of 755 kg and a velocity of 22.5 m/s. After the collision, the two cars stuck together. Considering the collision as perfectly inelastic. What will be the resulting velocity of the two cars together? Express your final answer in one decimal place rounded. *

Your answer

Submit

Clear form




1
Expert's answer
2021-11-07T19:29:08-0500
"m_1v_1+m_2v_2=(m_1+m_2)v_{res},""v_{res}=\dfrac{m_1v_1+m_2v_2}{m_1+m_2},""v_{res}=\dfrac{2500\ kg\times30.5\ \dfrac{m}{s}+755\ kg\times22.5\ \dfrac{m}{s}}{2500\ kg+755\ kg}=28.64\ \dfrac{m}{s}."